OpenStudy (anonymous):

y is directly proportional to x^2. It is known that y=10 for a particular value of x. Find the value of y when the value of x is halved

OpenStudy (amistre64):

doesnt y being directly proportional mean that: y = kx ? or in this case y=kx^2

OpenStudy (amistre64):

10 = kn^2 k = 10/n^2 y = 10/n^2 * (n/2)^2 y = 10n^2/4n^2 y = 10/4